OpenStudy (anonymous):

Given the sequence in the table below, determine the sigma notation of the sum for term 4 through term 15. n an 1 5 2 −10 3 20

OpenStudy (campbell_st):

well it looks like the term is \[5 \times (-2)^{n - 1}\] so the sigma notation is \[\sum_{n = a}^{b} 5 \times (-2)^{n - 1}\] you just need the a and b values...

OpenStudy (anonymous):

how do i do that @campbell_st

OpenStudy (anonymous):

@Melodysim

OpenStudy (anonymous):

Okay you dont have to use that @brynna

OpenStudy (campbell_st):

well it you look at \[5\times (-2)^{n - 1}\] when n = 1 the answer is \[5 \times (-2)^{1 - 1} = 5\] when n = 2 its \[5 \times (-2)^{2 - 1} = -10\] if you keep going you get the pattern so the sigma notation is an abbreviation for the sum of the terms... the question says start at 4 and finish the sum at 15 so its \[\sum_{n = 4}^{15} 5 \times (-2)^{n - 1}\]

OpenStudy (anonymous):

geometric sequence

OpenStudy (anonymous):

so that will be the formula.

OpenStudy (anonymous):

now a1 stands for first term

OpenStudy (anonymous):